# A differentiable Gillespie algorithm for simulating chemical kinetics, parameter estimation, and designing synthetic biological circuits

### v2

@@ Abstract
One prominent computational algorithm for understanding stochasticity in gene expression is the Gillespie algorithm (also known as the Stochastic Simulation Algorithm) [11, 12]. The Gillespie algorithm is an extremely efficient computational technique used to simulate the time evolution of a system in which events occur randomly and discretely [12]. Beyond gene expression, the Gillespie algorithm is widely employed across numerous disciplines to model stochastic systems characterized by discrete, randomly occurring events including epidemiology [13], ecology [14, 15], neuroscience [16, 17], and chemical kinetics [18, 19].
Here, we revisit the Gillespie algorithm in light of the recent progress in deep learning and differentiable programming by presenting a “fully-differentiable” variant of the Gillespie algorithm we dub the Differentiable Gillespie Algorithm (DGA). The DGA modifies the traditional Gillespie algorithm to take advantage of powerful automatic differentiation libraries (for example, PyTorch [20] and Jax [21]) and gradient-based optimization. The DGA allows us to quickly fit kinetic parameters to data and design discrete stochastic systems with a desired behavior. Our work is similar in spirit to other recent work that seeks to harness the power of differentiable programming to accelerate scientific simulations [22, 23, 24, 25, 26, 27]. The DGA’s use of differential programming tools also complements more specialized numerical methods designed for performing parameter sensitivity analysis on Gillespie simulations such as finite-difference methods [28, 29, 30], the likelihood ratio method [31, 32, 33] and pathwise derivative methods [34].
@@ I A differentiable approximation to the Gillespie Algorithm
Figure 1: Comparison between the exact Gillespie algorithm and the DGA for simulating chemical kinetics. (a) Example of kinetics with $N=3$ reactions with rates $r_{i}(i=1,2,3)$. (b) Visualization of approximations made in DGA. (c) Flow chart comparing exact and differentiable Gillespie simulations.
@@ I.3 Combining the DGA with gradient-based optimization
The goal of the optimization using the DGA is to find parameters $\bm{\theta}$ that minimize the loss. The basic workflow of a DGA-based optimization is shown in Fig. 2. One starts with an initial guess for the parameters $\bm{\theta}_{0}$. One then uses DGA algorithm to simulate the systems and calculate the gradient of the loss function $\nabla_{\theta}L(\bm{\theta})$. One then updates the parameters, moving in the direction of the gradient using gradient descent or more sophisticated second-order methods such as ADAM [37, 38], until one reaches a local minimum of the loss function.
@@ II.1 Two-state promoter model
Figure 4: Accuracy of the DGA in simulating the two-state promoter architecture in Fig. 3(a). Comparison between the DGA and exact simulations for (a) steady-state mRNA distribution, (b) moments of the steady-state mRNA distribution, and (c) the probability for the promoter to be in the “ON” or “OFF” state. (d) Ratio of the Jensen-Shannon divergence $\text{JSD}\left(p_{\text{DGA}}||p_{\text{exact}}^{\text{ss}}\right)$ between the differentiable Gillespie PDF $p_{\text{DGA}}$ and the exact steady-state PDF $p_{\text{exact}}^{\text{ss}}$, and the Shannon entropy $\text{H}(p^{\text{ss}}_{\text{exact}})$ of the exact steady-state PDF. In all of the plots, 2000 trajectories are used. The simulation time used in panels (a)-(c) is marked by blue ‘x’. Parameter values: $k^{\text{R}}_{\text{on}}=0.5$, $k^{\text{R}}_{\text{off}}=1.0$, $r=10$, $\gamma=1$, $2001/a=$ 200, and $201/b=$ 20.
@@ II.2 Characterizing errors due to approximations in the DGA
Next, we compare the accuracy of the DGA in simulating mRNA abundance distributions across a range of simulation times (see Fig. 4(d)). The accuracy is quantified by the ratio of the Jensen-Shannon divergence $\text{JSD}\left(p_{\text{DGA}}||p_{\text{exact}}^{\text{ss}}\right)$ between the differentiable Gillespie PDF $p_{\text{DGA}}$ and the exact steady-state PDF $p_{\text{exact}}^{\text{ss}}$, and the entropy $\text{H}(p^{\text{ss}}_{\text{exact}})$ of the exact steady-state PDF. For probability distributions $P$ and $Q$ over the same discrete space $\mathcal{X}$, the JSD and H are defined as:
@@
As expected, the $\frac{\text{JSD}}{\text{H}}$ ratio decreases with increasing simulation time, indicating convergence towards the steady-state distribution of the exact Gillespie simulation, which is obtained at a simulation time of $10^{4}$. The saturation of the $\frac{\text{JSD}}{\text{H}}$ ratio at approximately 0.003 for long simulation times is due to the finite values of $a^{-1}$ and $b^{-1}$. In percentage terms, this ratio represents a 0.3% divergence, meaning that the DGA’s approximation introduces only a 0.3% deviation from the exact distribution, relative to the total uncertainty (entropy) in the exact system.
Finally, the bar plot in Fig. 4(c) shows simulation results for the probability of the promoter being in the “OFF” and “ON” states as predicted by the DGA (in red) and the exact Gillespie simulation (in blue). The differentiable Gillespie over-estimates the probability of being in the “OFF” state and underestimates the probability of being in the “ON” state. Nonetheless, given the discrete nature of this system, the DGA does a reasonable job of matching the results of the exact simulations.
@@ III Parameter Estimation using the DGA
In many applications, one often wants to estimate kinetic parameters from experimental measurements of a stochastic system. For example, in the context of gene expression, biologists are often interested in understanding biophysical parameters such as the rate at which promoters switch between states or a transcription factor unbinds from DNA. However, estimating kinetic parameters in stochastic systems poses numerous challenges because the vast majority of methods for parameter estimation are designed with deterministic systems in mind. Moreover, it is often difficult to analytically calculate likelihood functions making it difficult to perform statistical inference. One attractive method for addressing these difficulties is to combine differentiable Gillespie simulations with gradient-based optimization methods. By choosing kinetic parameters that minimize the difference between simulations and experiments as measured by a loss function, one can quickly and efficiently estimate kinetic parameters and error bars.
@@ III.2 Confidence intervals and visualizing loss landscapes
One shortcoming of Eq. (13) is that it treats each parameter in isolation and ignores correlations between parameters. On a technical level, this is reflected in the observation that the confidence intervals only know about the diagonal elements of the full Hessian $\partial^{2}_{ij}L(\bm{\theta})$. This shortcoming is especially glaring when there are many sets of parameters that all optimize the loss function [42, 43]. As discuss below, this is often the case in many stochastic systems including the two-state promoter architecture in Fig. 3. For this reason, it is often useful to make two dimensional plots of the loss function $L(\bm{\theta})$. To do so, for each pair of parameters, we simply sample the parameters around their optimal value and forward simulate to calculate the loss function $L(\bm{\theta})$. We then use this simulations to create two-dimensional heat maps of the loss function, allowing us to identify “soft directions” in parameter space.
@@ III.4 Parameter estimation on experimental data
Figure 7: Fitting of experimental data from Ref. [35] using the DGA. (a) Comparison between theoretical predictions from the DGA (solid curves) and experimental values of mean and the Fano factor for the steady-state mRNA levels are represented by square markers, along with the error bars, for two different promoters, lacUD5 and 5DL1. Solid curves are generated by using DGA to estimate $\hat{r}$, $\hat{\gamma}$, and $\{\hat{k}^{\text{R}}_{\text{on}}\}$ and using this as input to exact analytical formulas. (b) Comparison between the inferred values of $\frac{\hat{r}}{\hat{\gamma}}$ using DGA with experimentally measured values of this parameter from Ref. [35].
@@
The results of this procedure are summarized in Fig. 7. We find that for the lacUD5 promoter $\hat{r}=90.25$, $\hat{\gamma}=6.20$ and that $\hat{k}^{\text{R}}_{\text{on}}$ varies from a minimum value of 0.18 to a maximum value of 99.0. For the 5DL1 promoters $\hat{r}=87.48$ and $\hat{\gamma}=9.80$ and $\hat{k}^{\text{R}}_{\text{on}}$ varies between 3.64 and 99.0. Recall that we have normalized all rates to the repressor unbinding rate $k^{\text{R}}_{\text{off}}=1$. These values indicate that mRNA transcription occurs much faster compared to the unbinding of the repressor, suggesting that once the promoter is in an active state, it produces mRNA rapidly. The relatively high mRNA degradation rates indicate a mechanism for fine-tuning gene expression levels, ensuring that mRNA does not persist too long in the cell, which could otherwise lead to prolonged expression even after promoter deactivation. The broad range of repressor binding rates shows that the system can adjust its sensitivity to repressor concentration, allowing for both tight repression and rapid activation depending on the cellular context.
@@
An appealing feature of [35] is that the authors performed independent experiments to directly measure the normalized transcription rate $r/\gamma$ (namely the ratio of the transcription rate and the mRNA degradation rate). This allows us to compare the DGA predictions for these parameters to ground truth measurements of kinetic parameters. As can be seen in Fig. 7(b), the predictions of the DGA agree remarkably well for both the lacUD5 and 5DL1 promoters.
@@ V Conclusion
In this paper, we introduced a fully differentiable variant of the Gillespie algorithm, the DGA. By integrating differentiable components into the traditional Gillespie algorithm, the DGA facilitates the use of gradient-based optimization techniques, such as gradient descent, for parameter estimation and network design. The ability to smoothly approximate the discrete operations of the traditional Gillespie algorithm with continuous functions facilitates the computation of gradients via backpropagation, a foundational technique in machine learning and has the potential to significantly expands the utility of stochastic simulations. Our work demonstrates the efficacy of the DGA through various applications, including parameter learning and the design of simple gene regulatory networks.

### v3

@@ Abstract
One prominent computational algorithm for understanding stochasticity in gene expression is the Gillespie algorithm, with its Direct Stochastic Simulation Algorithm variant being the most commonly used method [11, 12]. The Gillespie algorithm is an extremely efficient computational technique used to simulate the time evolution of a system in which events occur randomly and discretely [12]. Beyond gene expression, the Gillespie algorithm is widely employed across numerous disciplines to model stochastic systems characterized by discrete, randomly occurring events including epidemiology [13], ecology [14, 15], neuroscience [16, 17], and chemical kinetics [18, 19].
Here, we revisit the Gillespie algorithm in light of the recent progress in deep learning and differentiable programming by presenting a “fully-differentiable” variant of the Gillespie algorithm we dub the Differentiable Gillespie Algorithm (DGA). The DGA modifies the traditional Gillespie algorithm to take advantage of powerful automatic differentiation libraries (for example, PyTorch [20], Jax [21], and Julia [22]) and gradient-based optimization. The DGA allows us to quickly fit kinetic parameters to data and design discrete stochastic systems with a desired behavior. Our work is similar in spirit to other recent work that seeks to harness the power of differentiable programming to accelerate scientific simulations [23, 24, 25, 26, 27, 28, 29]. The DGA’s use of differential programming tools also complements more specialized numerical methods designed for performing parameter sensitivity analysis on Gillespie simulations such as finite-difference methods [30, 31, 32], the likelihood ratio method [33, 34, 35] and pathwise derivative methods [36].
@@ I A differentiable approximation to the Gillespie Algorithm
Figure 1: Comparison between the exact Gillespie algorithm and the DGA for simulating chemical kinetics. (a) Example of kinetics with $N=3$ reactions with rates $r_{i}(i=1,2,3)$. (b) Illustration of the DGA’s approximations: replacing the non-differentiable Heaviside and Kronecker delta functions with smooth sigmoid and Gaussian functions, respectively. (c) Flow chart comparing exact and differentiable Gillespie simulations.
@@ I.3 Combining the DGA with gradient-based optimization
The goal of the optimization using the DGA is to find parameters $\bm{\theta}$ that minimize the loss. The basic workflow of a DGA-based optimization is shown in Fig. 2. One starts with an initial guess for the parameters $\bm{\theta}_{0}$. One then uses DGA algorithm to simulate the systems and calculate the gradient of the loss function $\nabla_{\theta}L(\bm{\theta})$. One then updates the parameters, moving in the direction of the gradient using gradient descent or more advanced methods such as ADAM [39, 40], which uses adaptive estimates of the first and second moments of the gradients to speed up convergence to a local minimum of the loss function.
@@ II.1 Two-state promoter model
Figure 4: Accuracy of the DGA in simulating the two-state promoter architecture in Fig. 3(a). Comparison between the DGA and exact simulations for (a) steady-state mRNA distribution, (b) moments of the steady-state mRNA distribution, and (d) the probability for the promoter to be in the “ON” or “OFF” state. (c) Ratio of the Jensen-Shannon divergence $\text{JSD}\left(p_{\text{DGA}}||p_{\text{exact}}^{\text{ss}}\right)$ between the differentiable Gillespie PDF $p_{\text{DGA}}$ and the exact steady-state PDF $p_{\text{exact}}^{\text{ss}}$, and the Shannon entropy $\text{H}(p^{\text{ss}}_{\text{exact}})$ of the exact steady-state PDF. In all of the plots, 2000 trajectories are used. The simulation time used in panels (a),(b), and (d) is marked by blue ‘x’. Parameter values: $k^{\text{R}}_{\text{on}}=0.5$, $k^{\text{R}}_{\text{off}}=1.0$, $r=10$, $\gamma=1$, $2001/a=$ 200, and $201/b=$ 20.
@@ II.2 Characterizing errors due to approximations in the DGA
Next, we compare the accuracy of the DGA in simulating mRNA abundance distributions across a range of simulation times (see Fig. 4(c)). The accuracy is quantified by the ratio of the Jensen-Shannon divergence $\text{JSD}\left(p_{\text{DGA}}||p_{\text{exact}}^{\text{ss}}\right)$ between the differentiable Gillespie PDF $p_{\text{DGA}}$ and the exact steady-state PDF $p_{\text{exact}}^{\text{ss}}$, and the entropy $\text{H}(p^{\text{ss}}_{\text{exact}})$ of the exact steady-state PDF. For probability distributions $P$ and $Q$ over the same discrete space $\mathcal{X}$, the JSD and H are defined as:
@@
The ratio $\frac{\mathrm{JSD}}{H}$ normalizes divergence by entropy, enabling meaningful comparison across systems. As expected, the $\frac{\text{JSD}}{\text{H}}$ ratio decreases with increasing simulation time, indicating convergence towards the steady-state distribution of the exact Gillespie simulation. By “steady-state distribution”, we mean the long-term probability distribution of states that the exact Gillespie algorithm approaches after a simulation time of $10^{4}$. The saturation of the $\frac{\text{JSD}}{\text{H}}$ ratio at approximately 0.003 for long simulation times is due to the finite values of $a^{-1}$ and $b^{-1}$. In percentage terms, this ratio represents a 0.3% divergence, meaning that the DGA’s approximation introduces only a 0.3% deviation from the exact distribution, relative to the total uncertainty (entropy) in the exact system.
Finally, the bar plot in Fig. 4(d) shows simulation results for the probability of the promoter being in the “OFF” and “ON” states as predicted by the DGA (in red) and the exact Gillespie simulation (in blue). The differentiable Gillespie over-estimates the probability of being in the “OFF” state and underestimates the probability of being in the “ON” state. Nonetheless, given the discrete nature of this system, the DGA does a reasonable job of matching the results of the exact simulations.
@@ III Parameter Estimation using the DGA
In many applications, one often wants to estimate kinetic parameters from experimental measurements of a stochastic system [44, 45, 46, 47]. For example, in the context of gene expression, biologists are often interested in understanding biophysical parameters such as the rate at which promoters switch between states or a transcription factor unbinds from DNA. However, estimating kinetic parameters in stochastic systems poses numerous challenges because the vast majority of methods for parameter estimation are designed with deterministic systems in mind. Moreover, it is often difficult to analytically calculate likelihood functions making it difficult to perform statistical inference. One attractive method for addressing these difficulties is to combine differentiable Gillespie simulations with gradient-based optimization methods. By choosing kinetic parameters that minimize the difference between simulations and experiments as measured by a loss function, one can quickly and efficiently estimate kinetic parameters and error bars.
@@ III.2 Confidence intervals and visualizing loss landscapes
One shortcoming of Eq. (13) is that it treats each parameter in isolation and ignores correlations between parameters. On a technical level, this is reflected in the observation that the confidence intervals only know about the diagonal elements of the full Hessian $\partial^{2}_{ij}L(\bm{\theta})$. This shortcoming is especially glaring when there are many sets of parameters that all optimize the loss function [48, 49]. As discuss below, this is often the case in many stochastic systems including the two-state promoter architecture in Fig. 3. For this reason, it is often useful to make two dimensional plots of the loss function $L(\bm{\theta})$. To do so, for each pair of parameters, we simply sample the parameters around their optimal value and forward simulate to calculate the loss function $L(\bm{\theta})$. We then use this simulations to create two-dimensional heat maps of the loss function. This allows us to identify “soft directions” in parameter space, where the loss function $L(\bm{\theta})$ changes slowly, indicating weak sensitivity to specific parameter combinations.
@@ III.4 Parameter estimation on experimental data
Figure 7: Fitting of experimental data from Ref. [37] using the DGA. (a) Comparison between theoretical predictions from the DGA (solid curves) and experimental values of mean and the Fano factor for the steady-state mRNA levels are represented by square markers, along with the error bars, for two different promoters, lacUD5 and 5DL1. Solid curves are generated by using DGA to estimate $\hat{r}$, $\hat{\gamma}$, and $\{\hat{k}^{\text{R}}_{\text{on}}\}$ and using this as input to exact analytical formulas. (b) Comparison between the inferred values of $\frac{\hat{r}}{\hat{\gamma}}$ using DGA with experimentally measured values of this parameter from Ref. [37]. (c) Inferred $\hat{k}^{\text{R}}_{\text{on}}$ values as a function of the mean mRNA level.
@@
The results of this procedure are summarized in Fig. 7. We find that for the lacUD5 promoter $\hat{r}=90.25$, $\hat{\gamma}=6.20$ and that $\hat{k}^{\text{R}}_{\text{on}}$ varies from a minimum value of 0.18 to a maximum value of 99.0. For the 5DL1 promoters $\hat{r}=87.48$ and $\hat{\gamma}=9.80$ and $\hat{k}^{\text{R}}_{\text{on}}$ varies between 3.64 and 99.0. Recall that we have normalized all rates to the repressor unbinding rate $k^{\text{R}}_{\text{off}}=1$. These values indicate that mRNA transcription occurs much faster compared to the unbinding of the repressor, suggesting that once the promoter is in an active state, it produces mRNA rapidly. The relatively high mRNA degradation rates indicate a mechanism for fine-tuning gene expression levels, ensuring that mRNA does not persist too long in the cell, which could otherwise lead to prolonged expression even after promoter deactivation.
As expected, the repressor binding rates decrease with the mean mRNA level (see Fig. 7(c)). The broad range of repressor binding rates shows that the system can adjust its sensitivity to repressor concentration, allowing for both tight repression and rapid activation depending on the cellular context.
@@
An appealing feature of [37] is that the authors performed independent experiments to directly measure the normalized transcription rate $r/\gamma$ (namely the ratio of the transcription rate and the mRNA degradation rate). This allows us to compare the DGA predictions for these parameters to ground truth measurements of kinetic parameters. z. 7(b), the predictions of the DGA agree remarkably well for both the lacUD5 and 5DL1 promoters.
@@ V Conclusion
In this paper, we introduced a fully differentiable variant of the Gillespie algorithm, the DGA. By integrating differentiable components into the traditional Gillespie algorithm, the DGA facilitates the use of gradient-based optimization techniques, such as gradient descent, for parameter estimation and network design. The ability to smoothly approximate the discrete operations of the traditional Gillespie algorithm with continuous functions facilitates the computation of gradients via both forward-mode and reverse-mode automatic differentiation, foundational techniques in machine learning, and has the potential to significantly expand the utility of stochastic simulations. Our work demonstrates the efficacy of the DGA through various applications, including parameter learning and the design of simple gene regulatory networks.
@@
The DGA could also be extended to stochastic spatial systems by incorporating reaction-diffusion master equations or lattice-based models. Its differentiability may enable efficient optimization of spatially heterogeneous reaction parameters. However, such extensions may need to address computational scalability and stability in high-dimensional spaces, especially in processes such as diffusion-driven pattern formation or spatial gene regulation.